A fan blade is rotating with a constant angular acceleration of +8.5 rad/s2. At what point on the blade, as measured from the axis of rotation, does the magnitude of the tangential acceleration equal that of the acceleration due to gravity?
Answer #1

Consider a point at distance r from axis

Tangential acceleration, at = \alpha r = 8.5 r

For at =g = 9.8 m/s2

8.5 r = 9.8

Therefore,

r = 9.8÷8.5 = 1.15 m